OpenStudy (anonymous):

What is the sum of the geometric sequence –4, 24, –144, … if there are 6 terms? a. –159,963 b. 26,660 c. 26,661 d. 159,963 I think its b, am i right?

OpenStudy (campbell_st):

well you need the common ratio found by \[\frac{T _{2}}{T _{1}}=\frac{T _{3}}{T _{2}}...\] as this is an alternating sequence the common ratio is negative then use the formula \[s _{6}=\frac{-4(r^6 -1)}{r-1}\]

OpenStudy (anonymous):

yea thanks i did that but im jsut checking to see if someone else got that answer as well.

OpenStudy (campbell_st):

use wolfram if you just want to check

OpenStudy (anonymous):

ok thanks~!